OpenStudy (anonymous):

Which of the following can not be evaluated using the double angle formulas and explain why. a. 2 sin75 cos75 b. 1 - 2 sin^2 3pi/8 c. (2 tan 15)/(1 - tan^2 15) d. 2 cos^2 12 -1

OpenStudy (xapproachesinfinity):

2cos75sin75=sin(2*75)=sin(150) can be evaluated 1-2sin^2(3pi/8)=cos(2*3pi8)=cos (3pi/4) also can be evaluated follow the same reasoning tell you find the needed one

OpenStudy (anonymous):

Oh snap I didn't notice the 2 in the numerator for the last one, therefore it can't be evaluated. Thanks for that!

OpenStudy (xapproachesinfinity):

last one didn;t have any denominator so i don;t get why you mentioned numerator

OpenStudy (anonymous):

second last one*

OpenStudy (anonymous):

no wait hold up

OpenStudy (anonymous):

i was looking at the wrong formula, c can be tan30x and d can be cos 24x so im assuming the last one because we don't know any special triangles that have 24 of 12 in their angles?